OpenStudy (anonymous):

If a/6=b/10=c/4 then a+b+c/c=?

OpenStudy (raden):

let the value of a/6=b/10=c/4 = k so, a/6 = k ---> a=6k b/10 = k ---> b=10k c/4 = k ---> c = 4k therefore the value of (a+b+c)/c = (6k+10k+4k)/4k = 20k/4k = 5

OpenStudy (anonymous):

thanks a lot ur correct